### v1

@@ 5.2 Image Clustering
While our experiments in the previous section are all based on a supervised learning setting, either by making use of known templates or choosing some part of the dataset as training data, we now transition to an unsupervised setting. We again consider the three datasets from before, now with $100$ samples per class, and study unsupervised classification in the sense that we first perform a cluster analysis on a subset of the data and, thereon, classify the remaining part based on the corresponding cluster centres in feature space.
@@
class:   1    2    3   centre:   1    2    3     train/test:  /  1   /  2   /  3
Eucl. R-CDT $\mathrm{m}$NR-CDT
Refer to caption Refer to caption Refer to caption
$\mathrm{tv}$NR-CDT
@@
In all our numerical experiments, we use $850$ radii and $256$ angles for the Radon transform and $256$ interpolation points for the CDT. The results for the academic dataset are reported in Table 7. We observe that our $h_{\mathrm{}}$NR-CDT feature representations clearly outperform the Euclidean baseline and R-CDT embedding, yielding perfect RI and VI values on both the train and test set. Moreover, the cluster visualization in Fig. 7 indicates that the classes are indeed mapped to single points in $h_{\mathrm{}}$NR-CDT space, as predicted by our theory.
@@
In case of the polygon dataset we observe that $\mathrm{tv}$NR-CDT outperforms even the $\mathrm{m}$NR-CDT with a near perfect Rand index of $97\%$, see Table 8. In contrast to this, $\mathrm{m}$NR-CDT yields a Rand index of $86\%$, which is still significantly better than the Euclidean baseline and R-CDT. An inspection of the cluster visualization in Fig. 9 reveals that now the classes are not mapped to singletons in $h_{\mathrm{}}$NR-CDT space. This, however, is to be expected as the polygon dataset includes small non-affine perturbations. Nevertheless, the classes are mapped to well-separated sets, which is more pronounced for $\mathrm{tv}$NR-CDT than $\mathrm{m}$NR-CDT. This again indicates the potential of including the whole angular information instead of only considering the maximum.
class:   1    5    7    centre:   1    5    7      train/test:  /  1   /  5   /  7
Eucl. R-CDT
Refer to caption Refer to caption
$\mathrm{m}$NR-CDT $\mathrm{tv}$NR-CDT
Refer to caption Refer to caption
@@
Finally, for the LinMNIST dataset $\mathrm{tv}$NR-CDT and $\mathrm{m}$NR-CDT yield comparable results and perform significantly better than the Euclidean and R-CDT embedding, which are nearly on the same level, cf. Table 9. The cluster visualizations in Fig. 8 show that $\mathrm{m}$NR-CDT and $\mathrm{tv}$NR-CDT yield nicely separated clusters with larger distances in $\mathrm{tv}$NR-CDT space, which might explain the slightly better results for $\mathrm{tv}$NR-CDT.
@@ 5.3 Multi-dimensional Classification
5.3.1 3D Object Recognition
The first generalized $h_{\mathrm{}}$NR-CDT we study numerically is the multi-dimensional extension in § 4.1, which can be applied to the recognition of 3d objects. For first proof-of-concept experiments, we consider the following two datasets:
@@ 5.3.2 Classification on the 3D Rotation Group
In our last set of numerical experiments we leave the Euclidean setting and study a classification task on the 3d rotation group $\mathrm{SO}(3)$ as in § 4.3. To this end, we consider a synthetic dataset based on three empirical template measures, each supported on $1000$ rotation matrices. More precisely, the matrices for the first template are generated by the Matrix-Fisher distribution [50] around $I_{3}$ with concentration $\kappa=10$, the matrices for the second template are uniformly distributed on the subset of rotations around the $x$-$y$-equator, and the matrices for the third template are constructed by computing the QR decomposition of random Gaussian matrices in $\mathbb{R}^{3\times 3}$. Thereon, these template measures are transformed by applying random rotations to form the dataset, we refer to as rotation dataset, see Fig. 14 for visualization.
We again adapt the NT classification experiments from § 5.1.1 and label the transformed data based on the nearest template in $h_{\mathrm{}}$NR-CDT space with respect to the Euclidean norm. The $h_{\mathrm{}}$NR-CDTs of the entire dataset are depicted in Fig. 13, where the underlying CDTs are sampled on an equispaced grid consisting of $1000$ points in $(0,1)$, and show an almost perfect alignment within each class. The achieved classification accuracies are reported in Table 12. As expected, all employed $h_{\mathrm{}}$NR-CDT variants yield perfect results, whereas R-CDT cannot distinguish the different classes.
  class 1     class 2     class 3
$\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{d}}$NR-CDT
Refer to caption Refer to caption Refer to caption
2 0.633 0.933 0.933 0.433
4 0.400 1.000 1.000 0.733
8 0.333 1.000 1.000 0.900
16 0.367 1.000 1.000 0.967
32 0.367 1.000 1.000 1.000
64 0.400 1.000 1.000 1.000
@@
Figure 14: Illustration of the rotation dataset by visualizing the action of the underlying rotation matrices on a cone pointing towards the north pole, whose surface is coloured by a periodic colour bar to depict rotations around its z-axis.
@@ 6 Conclusion
In this paper, we introduced the novel generalized $h_{\mathrm{}}$NR-CDT for feature representation and proved linear separability of classes generated by affine transforms of given templates. This was validated by numerical experiments in two- and three-dimensional settings showing a significant increase in classification accuracy over the original R-CDT. Potential future directions include the investigation of $h_{\mathrm{}}$NR-CDT performance in combination with more advanced classification and clustering methods, like in [28, 52, 53], as well as a more in-depth numerical study in various real-world applications.

### v2

@@ 5.2 Image Clustering
While our experiments in the previous section are all based on a supervised learning setting, either by making use of known templates or choosing some part of the dataset as training data, we now transition to an unsupervised setting. We consider the first three datasets from before, now with $100$ samples per class, and study unsupervised classification in the sense that we first perform a cluster analysis on a subset of the data and, thereon, classify the remaining part based on the corresponding cluster centres in feature space.
@@
class:   1    2    3   centre:   1    2    3     train/test:  /  1   /  2   /  3
Eucl. R-CDT $\mathrm{m}$NR-CDT
Refer to caption Refer to caption Refer to caption
$\mathrm{tv}$NR-CDT
@@
In all our numerical experiments, we use $850$ radii and $256$ angles for the Radon transform and $256$ interpolation points for the CDT. The results for the academic dataset are reported in Table 9. We observe that our $h_{\mathrm{}}$NR-CDT feature representations clearly outperform the Euclidean baseline and R-CDT embedding, yielding perfect RI and VI values on both the train and test set. Moreover, the cluster visualization in Figure 8 indicates that the classes are indeed mapped to single points in $h_{\mathrm{}}$NR-CDT space, as predicted by our theory.
@@
In case of the polygon dataset we observe that $\mathrm{tv}$NR-CDT outperforms even the $\mathrm{m}$NR-CDT with a near perfect Rand index of $97\%$, see Table 10. In contrast to this, $\mathrm{m}$NR-CDT yields a Rand index of $86\%$, which is still significantly better than the Euclidean baseline and R-CDT. An inspection of the cluster visualization in Figure 9 reveals that now the classes are not mapped to singletons in $h_{\mathrm{}}$NR-CDT space. This, however, is to be expected as the polygon dataset includes small non-affine perturbations. Nevertheless, the classes are mapped to well-separated sets, which is more pronounced for $\mathrm{tv}$NR-CDT than $\mathrm{m}$NR-CDT. This again indicates the potential of including the whole angular information instead of only considering the maximum.
@@
Finally, for the LinMNIST dataset $\mathrm{tv}$NR-CDT and $\mathrm{m}$NR-CDT yield comparable results and perform significantly better than the Euclidean and R-CDT embedding, which are nearly on the same level, cf. Table 11. The cluster visualizations in Figure 10 show that $\mathrm{m}$NR-CDT and $\mathrm{tv}$NR-CDT yield nicely separated clusters with larger distances in $\mathrm{tv}$NR-CDT space, which might explain the slightly better results for $\mathrm{tv}$NR-CDT.
@@
class:   1    5    7    centre:   1    5    7      train/test:  /  1   /  5   /  7
Eucl. R-CDT
Refer to caption Refer to caption
$\mathrm{m}$NR-CDT $\mathrm{tv}$NR-CDT
Refer to caption Refer to caption
@@ 5.3 Multi-dimensional Classification
5.3.1 3D Object and Pose Recognition
The first generalized $h_{\mathrm{}}$NR-CDT we study numerically is the multi-dimensional extension in § 4.1, which can be applied to the recognition of 3d objects. For first proof-of-concept experiments, we consider the following three datasets:
@@
  3. c)
    FAUST dataset. The original ‘MPI FAUST Dataset’ [54] comprises real-world human scans of 10 persons in 30 poses. We use a subset of 10 poses from the registration data. Each sample is represented by a mesh with 6890 vertices, which we equip with the uniform distribution to obtain an empirical measure on $\mathbb{R}^{3}$. For illustration, four humans in four poses are visualized in Figure 13.
    human 1 human 3 human 5 human 7
    pose 1
    Refer to caption Refer to caption Refer to caption Refer to caption
    pose 2
    Refer to caption Refer to caption Refer to caption Refer to caption
    pose 3
    Refer to caption Refer to caption Refer to caption Refer to caption
    pose 4
    Refer to caption Refer to caption Refer to caption Refer to caption
    Figure 13: Visualization of a subset of the FAUST dataset: four humans in four poses.
In all cases, the data represents surfaces of 3d objects with positive volume so that the induced empirical measures are not supported on planes and, hence, are contained in $\mathcal{P}_{\mathrm{c}}^{*}(\mathbb{R}^{3})$.
@@
train Eucl. R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{c}}$NR-CDT
1 $0.658\pm 0.054$ $0.475\pm 0.060$ $\bf{0.706\pm 0.044}$ $0.672\pm 0.030$
3 $0.727\pm 0.035$ $0.575\pm 0.041$ $\bf{0.801\pm 0.050}$ $0.784\pm 0.042$
5 $0.757\pm 0.037$ $0.591\pm 0.040$ $\bf{0.838\pm 0.044}$ $0.828\pm 0.038$
Table 14: 1-NN classification accuracies for the FAUST dataset with $10$ samples per class and using 256 directions (in form of Fibonacci-like points on $\mathbb{S}_{2}$).
class:   1    2    3    4    5  centre:   1    2    3    4    5  train/test:  /  1   /  2   /  3   /  4   /  5
class:    6    7     8    9    10   centre: 6    7    8    9    10   train/test:/  6   /  7   /  8   /  9   /  10
Eucl. R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{c}}$NR-CDT
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: $10$-means cluster visualization for the FAUST dataset using a 3d PCA in the respective feature space.
Finally, we consider the FAUST dataset as benchmark for pose estimation. Since in this case no templates exist, we perform 1-NN classification, similar to the LinMNIST example in § 5.1. The resulting accuracies are shown in Table 14, where we use 256 directions on $\mathbb{S}^{2}$. We observe that $\mathrm{m}$NR-CDT performs best, followed by $h_{\mathrm{c}}$NR-CDT. Additionally, we perform a cluster analysis using $k$-means as in § 5.2. To this end, each class is split into a train and test subset, each consisting of five samples. The results are reported in Table 15 and the clusters are visualized in Figure 15. While $h_{\mathrm{c}}$NR-CDT shows the best Rand index on both the train and test sets, the clusters using $\mathrm{m}$NR-CDT visually appear more separated. We wish to stress that some poses are very similar so that a perfect classification/clustering cannot be expected.
Eucl. R-CDT $\mathrm{m}$NR-CDT $h_{\mathrm{c}}$NR-CDT
[…]
Table 15: Quality measures for $10$-means clustering of the FAUST dataset with $10$ samples per class, where 55 samples are used for training and the rest for testing.
@@ 5.3.2 Classification on the 3D Rotation Group
In our last set of numerical experiments we leave the Euclidean setting and study a classification task on the 3d rotation group $\mathrm{SO}(3)$ as in § 4.3. To this end, we consider a synthetic dataset based on three empirical template measures, each supported on $1000$ rotation matrices. More precisely, the matrices for the first template are generated by the Matrix-Fisher distribution [56] around $I_{3}$ with concentration $\kappa=10$, the matrices for the second template are uniformly distributed on the subset of rotations around the $x$-$y$-equator, and the matrices for the third template are constructed by computing the QR decomposition of random Gaussian matrices in $\mathbb{R}^{3\times 3}$. Thereon, these template measures are transformed by applying random rotations to form the dataset, we refer to as rotation dataset, see Figure 16 for visualization.
@@
Figure 16: Illustration of the rotation dataset by visualizing the action of the underlying rotation matrices on a cone pointing towards the north pole, whose surface is coloured by a periodic colour bar to depict rotations around its z-axis. For readability, we only show the first 20 out of 1000 components for each empirical measure.
  class 1     class 2     class 3
$\mathrm{m}$NR-CDT $h_{\mathrm{a}}$NR-CDT $h_{\mathrm{d}}$NR-CDT
Refer to caption Refer to caption Refer to caption
We again adapt the NT classification experiments from § 5.1.1 and label the transformed data based on the nearest template in $h_{\mathrm{}}$NR-CDT space with respect to the Euclidean norm. The $h_{\mathrm{}}$NR-CDTs of the entire dataset are depicted in Figure 17, where the underlying CDTs are sampled on an equispaced grid consisting of $1000$ points in $(0,1)$, and show an almost perfect alignment within each class. The achieved classification accuracies are reported in Table 16. As expected, all employed $h_{\mathrm{}}$NR-CDT variants yield perfect results, whereas R-CDT cannot distinguish the different classes.
2 0.633 0.933 0.933 0.433
4 0.400 1.000 1.000 0.733
8 0.333 1.000 1.000 0.900
16 0.367 1.000 1.000 0.967
32 0.367 1.000 1.000 1.000
64 0.400 1.000 1.000 1.000
@@ 6 Conclusion
In this paper, we introduced the novel generalized $h_{\mathrm{}}$NR-CDT for feature representation and proved linear separability of classes generated by affine transforms of given templates. This was validated by numerical experiments in two- and three-dimensional as well as non-Euclidean settings showing a significant increase in classification accuracy over the original R-CDT.
Potential future directions include the investigation of $h_{\mathrm{}}$NR-CDT performance in combination with more advanced classification and clustering methods, like in [32, 58, 59], as well as a more in-depth numerical study in various real-world applications. Moreover, the robustness of $h_{\mathrm{}}$NR-CDT to noise and outliers needs to be analyzed. For $\mathrm{m}$NR-CDT, this has already been done in our previous work [37] and we wish to adapt our techniques to the generalizations introduced here. Our theoretical results require the distinguishability of the embedded measure classes in feature space. While no collisions arose in our numerical experiments, a more descriptive characterization of the non-injectivity would still greatly enhance the application of our theory in practice.
At the current stage, our $h_{\mathrm{}}$NR-CDT is limited to probability measures. It may be extended to signed measures by adapting the signed CDT [60] and handling the positive and negative parts separately. Moreover, in our experiments, we always choose the uniform reference measure $\rho=u_{[0,1]}$. While our theory is independent of its choice, we expect that an application-dependent choice can even boost the $h_{\mathrm{}}$NR-CDT performance by, e.g., introducing a denser sampling of certain regions.